### v2

# KVCrush: Key Value Cache size-reduction using similarity in head-behaviour

Abstract
Key-value (KV) caching has emerged as a crucial optimization technique for accelerating inference in large language models (LLMs). By allowing the attention operation to scale linearly rather than quadratically with the total sequence length, KV caching significantly enhances generation throughput (h2o). However, due to large context lengths in the modern LLMs, the memory footprint of the KV is a huge bottleneck for model deployment directly impacting the model’s batch size, hindering its ability to deliver high-throughput (scissorhands). Existing research addresses this challenge using several techniques, such as discarding low-attention tokens, quantization, and matrix approximation which typically lead to a negative impact on the model accuracy.
In this paper, we propose KVCrush technology which can be combined with many KV compression technologies to improve the model accuracy at a much smaller memory. KVCrush provides an alternate representation scheme for key-value states, along with a low-overhead token pruning algorithm that accounts for the token distribution in the KV cache, which in turn allows for a smaller footprint while maintaining the accuracy of the model. Based on our results, KVCrush reduces LongBench longbench KV Cache size by with less than accuracy drop and achieves state-of-the-art average accuracy with minimal overhead, incurring less than total inference latency. KVCrush not only outperforms state-of-the-art importance-based token retention schemes in accuracy, but also integrates seamlessly with quantization, paging, and head-sharing techniques, requiring no retraining or architectural changes.
1 Introduction
Generative AI models, such as large language models (LLMs), have revolutionized the computing industry. These models boast an extensive number of parameters and consistently achieve state-of-the-art performance across various downstream tasks sun2019bert4rec memrec dosovitskiy2020image raffel2020exploring. However, the current trend of model size growth toward multi-trillion parameters isaev2023scaling — with models growing by one estimate gholami2024ai at a staggering rate of every 2 years— poses huge challenges for deploying them in practice. For instance, GPT-175B gpt3 requires 325GB of memory just to load the model weights. Additionally, not only do inferences for these models strain the platform compute and memory resources (both in terms of bandwidth and capacity), but typically, this is also coupled with strict latency requirements (in the order of tens of milliseconds) which further complicates the problem and poses a significant engineering challenge for efficiently delivering high inference throughput while maintaining low latency for user requests.
LLMs typically consist of stacked transformer decoder layers, with the self-attention module being a critical building block. This module weighs the importance of different tokens—capturing their contextual relationships—by attending over key–value token pairs at different positions within the input sequence. Nevertheless, self-attention matrix multiplication is compute intensive and has a quadratic computational complexity with respect to sequence length allyouneed which significantly impacts the inference throughput, especially for longer sequences survey.
KV caching has emerged as the de facto solution to this issue, converting the time complexity of token generation from quadratic to linear, albeit with increased memory overhead proportional to the context length. This mechanism is crucial for autoregressive tasks, where the model generates one output token at a time. In the decode phase, each token depends on the key and value tensors of all previous tokens (including the input tokens’ KV tensors computed at prefill and all new KV tensors computed until the current time step). To avoid recomputing these tensors for every token at each step, a KV cache stores them in memory, incrementally appending new entries to the cache as generation progresses.
The scalability of KV caching becomes a critical issue as models grow larger and more complex. The total memory used by a KV cache can be determined using Equation 1, where is the batch size, represents the number of layers in the model, represents the number of attention heads used, represents the dimensionality of the embeddings, is the length of context in tokens, is the number of bytes per value stored (e.g. 4B for FP32) and the factor is because two matrices for K and V are needed. As a result, for a given model, the KV cache size grows linearly with the maximum sequence length in the input context and the batch size, which, in practice, can result in an enormous KV cache size.
For instance, consider the OPT-175B model with its impressive 175 billion parameters, which consumes approximately 325 GB of memory. However, when using a batch size of 128 and a sequence length of only 8K, the KV cache requires around 4608 GB of memory. This is an order of magnitude (12X) larger than the model weights themselves. The issue of KV cache size has become increasingly prominent and is a significant cost factor in deploying large language models (LLMs). This is true, especially with the recent trend of LLM models that have been meticulously developed to scale up to handle extremely long context (for example, Google’s Gemini-pro-1.5 has shown to support a staggering 1M token context length Gemini
Several approaches have been proposed to mitigate the memory bottleneck associated with KV caching. Recent research have explored different optimizations of KV caching, including approaches such as low-rank decomposition of the KV cache(e.g., decomp) or pruning non-essential KV cache h2o; snapkv; scissorhands; fastgen, however, most of these techniques struggle to maintain the accuracy of the model at a smaller KV cache footprint when compared to the full model.
To address this challenge, we introduce KVCrush, a novel KV cache optimization that provides an alternate representation scheme for key-value states, along with a low-overhead token pruning algorithm that accounts for the token distribution in the KV cache. KVCrush, in turn, allows for a smaller footprint while maintaining the accuracy of the model and can be easily combined with many KV cache compression technologies. It is a modular add-on that augments existing methods by preserving contextual diversity, and requires no model retraining or architectural changes. Furthermore, KVCrush also works seamlessly with KV cache paging schemes (such as vLLM vllm) and mixed precision quantization mixed typically used in practical deployments. The technical contributions of this work can be summarized as follows:
Hardware-Efficient Alternative Representation of Tokens: We leverage attention score patterns across heads to generate a binary feature vector for each token. This binary alternative representation is much smaller than the original key and value vectors, yet it preserves enough semantic information, to convey token importance and similarities, which we use to prune or retain the tokens very efficiently.
Accuracy-aware Low-Overhead KV Optimization Algorithm: Our algorithm leverages the alternative binary representation of tokens and an anchor point to bucketize tokens into representative groups, using very efficient low-overhead distance calculations that scale linearly with the number of tokens. This approach ensures that different token groups are represented in the reduced KV memory footprint, thereby maintaining the inference accuracy of the model.
Practical Efficiency and Integration: We show that KVCrush can be efficiently combined with token eviction-based KV cache compression methods. It achieves state-of-the-art average accuracy on LongBench, with less than inference latency overhead, and offers up to cache reduction with under accuracy drop compared to the full cache.
2 Related Work
Techniques to reduce the size of KV Caches have received substantial research attention in recent years. Broadly speaking, we can divide these into three categories:
2.1 Quantizing KV Cache
Quantization is the process of reducing the precision of a model’s parameters and activations to lower bit-widths to save memory and computational resources. Quantization can be categorized into post-training quantization (PTQ) and quantization-aware training (QAT), with PTQ often preferred for large language models due to its lower resource requirements. xiao2023smoothquant, liu2024llmqat, and sheng2023flexgen demonstrated that quantizing queries, keys, and values to INT8 enables efficient attention operations. However, these methods apply quantization uniformly across all KV cache tokens, and thus will typically negatively impact the model generation accuracy.
On the other hand, research work such as mq1, mq2, mq3, mikv employ mixed precision quantization to allocate different number of bits to various model components or tensors, and thus enable a more efficient compression. This methods leverage the insight that different parts of a model exhibit varying levels of sensitivity to quantization.
As we discuss later in detail, KVCrush operates directly on the KV cache to determine which tokens to retain or evict. Since it does not alter the cache format itself, it is naturally complementary to quantization techniques. In fact, KVCrush can work seamlessly with mixed-precision KV caches by assigning different bitwidths to retained tokens based on their importance.
2.2 Sharing KV Cache
Multi-Query Attention (MQA) and Grouped Query Attention (GQA) are techniques developed to address the memory footprint issues LLMs allowing to share KV caches across heads. MQA, introduced by (mqa), reduces memory usage by sharing key and value representations across all attention heads, which enhances memory efficiency and inference speed, at the expense of generation quality. GQA, proposed by (gqa), extends this concept by grouping query heads to share KVs, balancing memory reduction with better performance retention. However, GQA involves higher training costs and complexity. Both methods offer trade-offs between memory efficiency and model performance, with MQA favoring memory savings and GQA providing a more balanced approach.
Intuitively, KVCrush is completely orthogonal to these KV cache sharing schemes. It operates directly on the deployed cache, regardless of whether key and value states are shared or grouped across attention heads.
2.3 Evicting inconsequential Keys and Values
This category is the closest related work to KVCrush. In this research category different methods aim to prune key-value (KV) pairs from cache after input processing, aiming to enhance decoding efficiency. By evicting tokens out of KV cache, memory consumption is reduced, facilitating support for larger batch sizes and longer context windows.
Different strategies for KV pruning and selectively dropping tokens from the KV cache have been proposed in recent research work, For example, in StreamLLM streamingllm, only the most recent tokens and attention sinks (first few tokens) are retained. H2O h2o and Scissorhands scissorhands utilize attention-based metrics to determine eviction, with H2O summing attention weights and Scissorhands considering the frequency of attention surpassing a threshold. FastGen fastgen combines these methods with heuristics for special tokens. SnapKV snapkv similarly depends on attention weights to prune the KV cache, but addresses the efficiency during prefill with long context, and due to the complexity of computing attention score for the whole context, it limits the observation window to the final tokens of the input prompt, thus reducing the complexity from to , where is the context length, while using max pooling to retain neighboring KVs. PyramidKV pyramidkv builds on SnapKV by configuring variable eviction rates across layers, with more aggressive pruning in later layers where attention is less evenly distributed. DynamicKV dynamickv adaptively resizes KV caches across layers based on task type. LaCache lacache introduces a ladder-shaped caching pattern that prioritizes temporal coverage across layers.
The common drawback across these existing methods is that model accuracy drops as the KV compression ratio increases. These works focus solely on selecting a set of pivotal tokens (using varying algorithms) and evict the rest, which reduces KV cache size but at the cost of generation quality. KVCrush addresses this by explicitly representing groups of evicted tokens via a small set of representative tokens, preserving context diversity and mitigating accuracy loss under high compression. It ensures that different token groups are reflected in the reduced KV memory footprint, thereby sustaining accuracy. As a result, KVCrush can be combined with many existing methods to remedy accuracy drop at the same KV cache budget. We present results in Section 4, combining KVCrush with state-of-the-art techniques and demonstrating improved performance at equal compression ratios.
| Benchmarks | LongBench | lm-eval-harness | |||
| Experiments | Section 4.2, Section 4.3.2 | Section 4.3.1, Section 4.3.3, Section 4.3.4, Section 4.3.5 | |||
| Datasets |
GSM8K and XSUM | |||
| Models | Mistral-7B-Instruct-v0.2, Meta-Llama-3-8B-Instruct | Phi-3-mini-4k-instruct, Meta-Llama-3-8B-Instruct, Llama-2-7b-chat-hf | |||
| Baselines | FullKV, H2O, SnapKV, PyramidKV | H2O | |||
| Paging Mode | Token Level, Chunk Level | Page Level (Page Size = 32) | |||
| Total Cache Budget (tokens) | 2048 | 672 = 32(initial)+512(middle)+128(recent) | |||
| Cache Budget Partitioning | 25% KVCrush + 75% Baseline | 25% KVCrush + 75% Baseline | |||
| Cuda Version | 12.2 | 12.2 | |||
| Pytorch Version | 2.4.1+cu121 | 2.4.1+cu121 |
| Cache Budget: 512 | Phi-3-mini-4k-instruct | Meta-Llama-3-8B-Instruct | Llama-2-7b-chat-hf | ||||
| hh-cl | Strict | Flexible | Strict | Flexible | Strict | Flexible | |
| H2O | 512-0 | 70.7 | 79.3 | 74.9 | 74.7 | 0.209 | 0.225 |
| kvcrush.random | 128-384 | 75.4 | 80.9 | 76.2 | 76.2 | 0.211 | 0.226 |
| kvcrush.mean | 128-384 | 75.2 | 80.6 | 76.5 | 76.4 | 0.21 | 0.229 |
| kvcrush.alternate | 128-384 | 74.6 | 80.9 | 75.7 | 75.6 | 0.212 | 0.227 |
3 KVCrush
The basic flow of KVCrush is shown in Figure 1. Given a certain total budget for the KV Cache, this is split into two smaller portions, represents the cache budget available to store the set of pivotal tokens. This set is determined based on the specifics of KV cache compression algorithm used (e.g. H20, SnapKV, PyramidKV, etc.). While represents the cache budget available to store along a set of representative tokens, these act as proxies of the evicted tokens, and are selected based on low-overhead grouping algorithm (discussed in Section 3.2) to ensure better representation of the whole context.
3.1 KVCrush Alternative Token Representation
In the KV cache tokens are represented as floating-point vectors of size each. is the size of embedding length and is typically not small. For example, for GPT-3 gpt3 D is 2048 dimensions while for LLaMA-65B llama it is set at 4096 dimensions. As previously mentioned, KVCrush will try to group tokens in order to select representative tokens from each group. As a result, it is essential to minimize the overhead of the grouping algorithm. The running time of any clustering algorithm in dimensions will be proportional to the value of .
To minimize this overhead, KVCrush tries to have an alternative representation of each token with a vector of length . The main challenge is how to find this smaller representation and yet still preserve sufficient semantic information of the tokens to be able to differentiate and group them based on their contextual relationship.
KVCrush draws some insights of alternative token representation from the FastGen research fastgen. The authors in this work demonstrated that distinct attention heads exhibit unique structural patterns and attend to different tokens differently. For example, some attention heads will attend more to special tokens, others will attend more to locality among tokens, others will attend more to punctuation, etc.
Building on these findings we can deduce that the attention score of a certain token across heads is a vector of length and will represent good semantic information about the contextual properties of that token. Keeping in mind that , where for example, is 96, 128 for GPT-3 gpt3 and LLaMA-65B llama, respectively.
To reduce the grouping overhead even more, KVCrush takes the alternative token representation one step further. Instead of using a floating point vector of size each, it converts that into a binary vector of size replacing the floating point with a single bit. The main insight is that, for a given head and attention score based token eviction algorithm, the binary decision of whether to retain or evict a token encodes important contextual signals (e.g., attention score, recency), which we use to build token-level semantic fingerprints.
Algorithm 1 depicts how KVCrush generates a binary representation of size for each input token. The process can be summarized in the following steps:
Compute the attention weight matrix: For each head , compute the attention weight matrix using the query and key projections. Note that this attention computation is already performed during inference, and KVCrush leverages these values without recomputation.
Apply a per-head threshold: For each token and head , compute the row-wise normalized attention score . A token is marked with a bit value of 1 if , and 0 otherwise. The threshold is chosen such that it retains a target fraction of tokens for each head, based on desired compression ratio.
Form the binary representation: For each token , collate the bit values across all heads into a binary vector , which captures its cross-head importance signature.
3.2 KVCrush Token Grouping and Cache Eviction
As previously mentioned, KVCrush will select tokens that act as proxies of the evicted tokens based to keep in the KV Cache. The selection is based on a low-overhead grouping algorithm built on top of the alternative binary representation of tokens. Algorithm 2 depicts the main steps of this weak clustering algorithm to form token groups and can be summarized as follows:
The clustering algorithm takes a set of tokens each represented with a binary vector of length as input, it selects tokens where vectors as output
An arbitrary anchor point is first selected (ideally, the anchor point should be centrally placed to all vectors in an dimension space). In our experiments we present results for 3 different (low-overhead) selection of anchor points (random, alternate and mean).
For each of the input vectors, hamming distance is computed with respect to the anchor point. The binary representation enables computing Hamming distances on-the-fly using lightweight bitwise operations, making the grouping process extremely efficient compared to standard clustering.
Each token is then assigned to one of the buckets. The selected bucket is the one with the lowest hamming distance.
After the bucketization of all tokens into their corresponding buckets, one representative vector nearest to the centroid of each bucket is selected, and the remaining tokens are dropped. The retained tokens, along with the tokens preserved by the KV compression algorithm, form the final tokens of the KV cache.
After assigning all tokens to their respective buckets, one representative vector is selected from each by finding the token whose binary representation is closest to that bucket’s centroid. All the other () tokens are dropped. Those retained tokens, in addition to the tokens retained by the KV compression algorithm, represent the final tokens of the KV cache.
It should be mentioned however, that we described KVCrush grouping algorithm using Hamming distance computations with respect to ONE anchor point, and thus, token pruning only requires Hamming distance comparisons instead of using distance computations required by standard clustering algorithms. In Section 4.3.3 we present results for using a higher overhead clustering algorithm on accuracy and latency of KVCrush.
4 Experiments
We organize our experiments to evaluate KVCrush along four dimensions: accuracy and latency trade-offs under constrained cache budgets; ablation studies on anchor selection, budget partitioning, and clustering methods (including comparisons with k-means); integration with existing KV eviction methods such as H2O, SnapKV, and PyramidKV; and performance in paged KV cache settings.
4.1 Experimental Setup
We summarize experimental setup used for Evaluation in table 1.
4.1.1 Baselines
For baseline comparison, we evaluate KVCrush alongside four methods: FullKV uses the complete KV cache with zero eviction. H2O (h2o) determines eviction based on the cumulative attention weights, retaining the top- tokens per layer. SnapKV (snapkv) restricts the observation window to the final tokens of the prompt while employing max pooling to preserve neighboring KVs. PyramidKV (pyramidkv) extends SnapKV by applying variable eviction rates across layers, pruning more aggressively in later layers where attention distributions are less uniform.
4.1.2 Datasets and Models
We evaluate KVCrush on 16 diverse LongBench datasets and the GSM-8K and XSUM tasks from lm-eval-harness, using LLaMa-3-8B-Instruct, Mistral-7B-Instruct-v0.2, and Phi-3-mini-4k-instruct models. Details of datasets, models, and input lengths are summarized in Table 1.
4.2 End-to-End Evaluation
4.2.1 Accuracy
| Single Document QA | Multi Document QA | Summarization | Few Shot Learning | Synthetic | Code | |||||||||||
narrativeqa |
qasper |
multifieldqa_en |
hotpotqa |
2wikimqa |
musique |
gov_report |
qmsum |
multi_news |
trec |
triviaqa |
samsum |
psg_count |
psg_ret_en |
repobench-p |
Average |
| 18409 | 3619 | 4559 | 9151 | 4887 | 11214 | 8734 | 10614 | 2113 | 5177 | 8209 | 6258 | 11141 | 9289 | 4206 | 7839 | |
| Compression (x) | 9.0 | 1.8 | 2.2 | 4.5 | 2.4 | 5.5 | 4.3 | 5.2 | 1.0 | 2.5 | 4.0 | 3.1 | 5.4 | 4.5 | 2.1 | 4 |
| Mistral-7B-Instruct-v0.2 (Cache Budget: 2048) | ||||||||||||||||
| FullKV | 26.85 | 33.06 | 49.44 | 43.02 | 27.33 | 18.78 | 32.87 | 24.21 | 27.05 | 71.00 | 86.23 | 42.90 | 2.75 | 86.98 | 54.41 | 41.79 |
| H2O | 25.10 | 30.67 | 48.29 | 40.89 | 25.98 | 15.57 | 28.06 | 23.48 | 26.78 | 60.50 | 86.33 | 42.48 | 2.57 | 82.73 | 52.92 | 39.49 |
| H2O+KVCrush | 25.77 | 30.78 | 48.47 | 41.29 | 26.33 | 16.42 | 28.87 | 23.62 | 26.83 | 65.42 | 86.26 | 42.93 | 2.67 | 83.44 | 53.21 | 40.15 |
| SnapKV | 26.15 | 32.38 | 49.54 | 41.66 | 27.54 | 19.43 | 29.58 | 23.83 | 26.70 | 71.00 | 86.31 | 43.07 | 2.89 | 85.89 | 53.87 | 41.32 |
| SnapKV+KVCrush | 26.20 | 32.09 | 49.86 | 41.86 | 28.33 | 18.85 | 29.62 | 23.92 | 26.95 | 71.00 | 86.20 | 42.99 | 2.88 | 86.06 | 54.13 | 41.40 |
| PyramidKV | 25.82 | 31.67 | 49.20 | 41.19 | 27.01 | 19.37 | 29.15 | 23.89 | 26.72 | 71.00 | 86.28 | 43.24 | 2.73 | 85.06 | 53.57 | 41.06 |
| PyramidKV+KVCrush | 26.05 | 31.99 | 49.30 | 41.37 | 27.12 | 18.79 | 29.15 | 24.22 | 27.05 | 71.00 | 86.25 | 43.01 | 2.77 | 85.23 | 53.87 | 41.14 |
| KVCrush* | 26.20 | 32.09 | 49.86 | 41.86 | 28.33 | 18.85 | 29.62 | 24.22 | 27.05 | 71.00 | 86.26 | 43.01 | 2.88 | 86.06 | 54.13 | 41.43 |
| LLaMa-3-8B-Instruct (Cache Budget: 2048) | ||||||||||||||||
| FullKV | 25.56 | 31.95 | 39.71 | 43.56 | 35.63 | 21.18 | 28.58 | 23.27 | 26.75 | 74.00 | 90.48 | 42.30 | 4.80 | 69.25 | 53.92 | 40.73 |
| H2O | 25.42 | 26.43 | 38.87 | 42.82 | 32.91 | 20.02 | 25.09 | 23.26 | 26.11 | 58.50 | 90.56 | 41.57 | 5.20 | 69.50 | 54.10 | 38.69 |
| H2O+KVCrush | 25.48 | 27.01 | 39.02 | 43.20 | 33.23 | 20.78 | 25.67 | 23.17 | 26.33 | 62.00 | 90.48 | 41.87 | 5.23 | 69.50 | 54.32 | 39.15 |
| SnapKV | 25.70 | 29.96 | 38.93 | 43.90 | 35.05 | 20.44 | 26.89 | 23.43 | 26.17 | 74.00 | 90.56 | 41.96 | 5.54 | 69.25 | 56.16 | 40.53 |
| SnapKV+KVCrush | 25.62 | 29.48 | 39.56 | 44.05 | 36.20 | 20.93 | 26.24 | 23.16 | 26.32 | 74.00 | 90.54 | 42.17 | 5.83 | 69.25 | 55.48 | 40.59 |
| PyramidKV | 25.53 | 29.89 | 38.67 | 43.90 | 35.04 | 21.60 | 26.80 | 23.51 | 26.37 | 73.50 | 90.56 | 42.21 | 5.08 | 69.25 | 55.36 | 40.48 |
| PyramidKV+KVCrush | 25.57 | 29.48 | 38.97 | 44.03 | 35.75 | 21.62 | 26.21 | 23.28 | 26.52 | 74.00 | 90.56 | 42.24 | 5.17 | 69.50 | 55.49 | 40.56 |
| KVCrush* | 25.62 | 29.48 | 39.56 | 44.05 | 36.20 | 21.62 | 26.24 | 23.28 | 26.52 | 74.00 | 90.56 | 42.24 | 5.83 | 69.50 | 55.49 | 40.68 |
The accuracy of KVCrush, in comparison with baseline methods, is detailed in Table 3. For this evaluation, a cache budget of 2048 was used. KVCrush employs a static cache budget partitioning scheme, wherein of the total budget (i.e., 512) is allocated to KVCrush for selecting representative tokens. The remaining is distributed to the accompanying method (H2O, SnapKV, or PyramidKV) for selecting high-attention tokens. The key insights are as follows:
KVCrush when coupled with the best performing KV compression method (between H2O, SnapKV and PyramidKV) achieves the best accuracy for most of the datasets on both LLaMa-3-8B-Instruct and Mistral-7B-Instruct-v0.2 models.
Using Mistral-7B-Instruct-v0.2, it is the only method to achieve iso-accuracy w.r.t. FullKV baseline on qmsum and multi-news datasets.
KVCrush achieves the best average accuracy on both LLaMa-3-8B-Instruct and Mistral-7B-Instruct-v0.2 models.
4.2.2 Inference Latency Overhead
fig:latency_breakdown
We evaluate the runtime latency of different KV cache strategies using a LongBench microbenchmark. The experiments were conducted on a system equipped with an Intel® Xeon® Platinum 8470 processor, simulating realistic CPU-bound server inference settings. The cache budget and associated compression parameters are detailed in Table 1.
Figure LABEL:fig:latency_breakdown shows the normalized latency breakdown across three components: memory access, compute, and pruning overhead. We observe that while the full cache baseline incurs the highest memory access latency, both H2O and H2O+KVCrush significantly reduce this cost by reducing the KV cache size by . Importantly, the additional pruning overhead introduced by KVCrush remains minimal—only higher than H2O—while achieving higher accuracy. This demonstrates that KVCrush can be deployed with negligible runtime cost while improving performance.
4.3 Ablation Study
4.3.1 Selecting Anchor Points
KVCrush employs an anchor point within the same binary space (a binary vector with a length equal to the number of heads) to form clusters of tokens based on their hamming distance to this anchor. As shown in Table 2, KVCrush outperforms the H2O baseline across all anchor-point strategies, including simple choices such as random, mean, and alternating 0–1 vectors. The 512-token budget shown is illustrative; similar trends are observed for other cache budgets.
4.3.2 Cache budget partitioning
As discussed in Section 3, the cache budget allocated for the KV Cache is divided into two distinct segments: , utilized by the baseline method to identify tokens with high attention weights, and , used by KVCrush to select representative tokens.
Figure 5 illustrates the impact of varying the percentage of the budget allocated to KVCrush on overall accuracy. For certain workloads, such as narrativeqa, the optimal cache budget can reach up to of the total budget. In contrast, for other workloads, the optimal budget typically ranges between and of the available budget. In Section 4.2.1, we compared our method with baseline approaches using a static cache budget partitioning scheme, allocating a fixed of the total budget to KVCrush. This approach could potentially be enhanced by dynamically partitioning the budget based on attention weights distribution, which we plan to explore in future work.
fig:kmeans
fig:kvcrush_int
4.3.3 KMeans vs KVCrush
Traditional clustering algorithms exhibit inefficiencies in KV Cache compression due to two primary factors. Firstly, the large size of input vectors poses a challenge. For instance, in Llama3-8B model, each input key (or value) vector comprises 128 FP16 values. By employing kvcrush.representation, a binary vector of length 32 (corresponding to the number of heads) is generated to represent each token (key-value pair). This approach not only reduces the data size by a factor of 64 but also enables faster distance computations using Hamming distances. Secondly, the time complexity for clustering algorithms is substantial. Utilizing kvcrush.grouping, the selection of values to be retained in the cache is achieved in time, in contrast to the time complexity of k-means clustering.
In Figure LABEL:fig:kmeans, we compare KVCrush and KMeans in terms of accuracy and inference latency using the GSM8K dataset. While KMeans achieves slightly higher accuracy than KVCrush (Figure LABEL:fig:kmeans_acc), it incurs approximately additional overhead in the inference latency pipeline. In contrast, KVCrush provides a reasonable accuracy improvement over the pure H2O baseline, with an insignificant overhead of less than .
4.3.4 Accuracy implications of integrating KVCrush with other methods
We assess the impact on accuracy when integrating KVCrush with other methods, utilizing cumulative attention-weights as the importance metric. Figure LABEL:fig:kvcrush_int demonstrates the accuracy improvement provided by KVCrush to the baseline methods for the 2wikimqa dataset. These baseline methods will prioritize selecting important tokens. As illustrated in Figure 1, KVCrush not only seamlessly integrates with these, but also improves upon their accuracies by selecting representative tokens (or pages) based on the head behavior. As shown in figure, KVCrush also enhances the accuracy of the baseline methods even when they operate at chunks instead of individual tokens.
fig:kvcrush_paged
4.3.5 Evaluation of KVCrush in Paged KV settings
To evaluate KVCrush in paged KV settings, we utilized an H2O baseline that aggregates the row-wise sum of attention weights at the page level to evict low-importance pages. For KVCrush, the binary vector of a page is formed by concatenating the binary head vectors of all tokens within that page. Figure LABEL:fig:kvcrush_paged shows that Paged-KVCrush outperforms paged-H2O on both GSM8K and XSUM datasets.
5 Conclusion and Future Work
In this work, we presented an alternative approach to represent LLM tokens during inference. We showed that a compact representation paired with KVCrush compression algorithm leads to a substantial reduction in KV cache memory footprint. We demonstrated how we can use this method to achieve a memory-efficient LLM inference pipeline without compromising the quality of the generated tokens. In future work, we intend to investigate dynamic cache budget allocation and develop a more refined multi-anchoring approach.